# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@ 2.3 Specification property
Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments using specification-like property.
Definition 2.4 (Specification property).
One says that a t.d.s. $(X,f)$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
$d(f^{p+a_{j}}x,f^{p}x_{j})<\epsilon$
@@
In other words, that point $x$ $\epsilon$-shadows the $k$ pieces of orbits
@@
A point $x\in X$ is almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, it holds that $f^{k}(x)\in U$ for some $n\leq k\leq n+N$. Let $AP(X)$ denote the set of all almost periodic points of $X$.
The following result, which was proved for systems with $g$-almost product property [HTZ23, Proposition 2.11], gives a characterization of the structure of $AP(X)$.
@@ 3 Proofs of main results
This section aims to prove the main results of this paper. Essentially, the difficult part is the estimation of lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction. Then such set is suitable for the purpose of the computation of the lower bound of metric mean dimensions of saturated sets.
@@ 3.1 Packing metric mean dimension of saturated sets
We first prove Theorem 1.1. To arrive this goal, we divide the proof of the Theorem 1.1 into two parts.
@@ 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
We adapt a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
@@
[Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
@@
This essentially follows the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 9.10]). ∎
For $K\subset M_{f}(X)$, we set $G^{K}:=\{x\in X:V_{f}(x)\subset K\}$. Obviously, we have the fact $G_{K}\subset G^{K}$.
@@
It suffices to show $(1)$, since one can take $K=\{\mu\}$ as in $(1)$ and use the fact $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
For $n\in\mathbb{N}$, $\delta>0$, we put
$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)~{}\text{for infinitely many }n>k\right\},$
@@
and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$. Fix $x\in G^{K}$. Then, we can choose a subsequence $\{n_{i}\}$ such that $\mathcal{E}_{n_{i}}(x)\to\mu$ for some $\mu\in K$. This implies for every $\delta>0$, there exists a positive integer $k$ such that $\mathcal{E}_{n_{i}}(x)\in B(K,\delta)$ for sufficiently large $n_{i}$ greater than $k$. Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ holds for every $\delta$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon),$
which implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon).$
@@
Take a strictly decreasing positive sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
$S(K,\delta_{k},n_{k},\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$, and set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x)\in B(K,\delta_{k}).$
Choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed convex set. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This completes the proof. ∎
Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of $G_{K}$:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,G_{K},\epsilon)}{|\log\epsilon|}\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.1)
@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
For any small enough $\epsilon^{\ast}>0$, we shall prove that
@@
One may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$. Then take $\mu^{\ast}\in K$ such that
@@
Since $K$ is connected and compact, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
@@
Now the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively. Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall write $\delta_{k}:=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
@@
For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$ which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2. Then according to specification property, there exists $a_{k}\in X$ such that
@@
Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$, $N_{(k,0)}=1$.
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation of $\mu^{*}$ derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
@@
For each $k\in\mathbb{N}$, there is a measure $\nu_{k}\in M_{f}(X)$ satisfying
@@
$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta.$ (3.4)
$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$.
By the ergodicity of $\nu_{i}^{k}$, there exists $N\in\mathbb{N}$ such that the set
@@
has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3 one has
@@
$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer holds for all $1\leq i\leq j_{k}$, and
@@
$(2)$ there exists a subset $E_{k,i}$ which is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
@@
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniform continuous.
For any distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$,
@@
Then $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set and
@@
The last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and formulas (3.3), (3.5).
For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$ taking values in $[0,1]$,
@@
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Recall we have
$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k}.$
By the construction, one has
@@
Until now, using the specification property we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the remaining cases for $1\leq j<Q_{k}$.
Since the system has specification property, $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$. Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist a point $x_{k,j}$ and $n_{(k,j)}\in\mathbb{N}$ such that
@@
Step $1$. Construction of intermediate sets $\mathcal{S}_{(k,j)}$.
@@
For $(k,j)\in\mathcal{A}$ with $j=0$, set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.

### v3

@@ 2.3 Specification property
Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as specification property.
Definition 2.4.
A t.d.s. $(X,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
$d(f^{p}x_{j},f^{p+a_{j}}x)<\epsilon$
@@
In other words, the point $x$ $\epsilon$-shadows the $k$ pieces of finite orbits of $x_{j}$:
@@
A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$. Let $AP(X)$ denote the set of all almost periodic points of $X$. The following statement, proved in [HTZ23, Proposition 2.11] for systems with $g$-almost product property, gives a characterization of the structure of $AP(X)$.
@@ 3 Proofs of main results
This section proves the main results of this paper. The challenging aspect lies in the estimation of the lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We will construct a Moran fractal set, which is geometrically constructed and is contained in $G_{K}^{\omega,U}$. Consequently, this set is well-suited for computing the lower bound of the metric mean dimensions of saturated sets.
@@ 3.1 Packing metric mean dimension of saturated sets
We first prove Theorem 1.1. To arrive at this goal, we divide its proof into two parts.
@@ 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
We employ a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to derive the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$.
@@
[Y23, Lemma 3.1] Let $(X,f)$ be a t.d.s. and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
@@
This essentially adheres to the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
Recall $G_{K}=\{x\in X:V_{f}(x)=K\}$. For $K\subset M_{f}(X)$, we set
$G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
Obviously, we have $G_{K}\subset G^{K}$.
@@
It suffices to show $(1)$ since one can take $K=\{\mu\}$ as in $(1)$ and use $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
For $n\in\mathbb{N}$ and $\delta>0$, we put
$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~{}\forall n\geq k\right\},$
@@
and
$\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon),$
where $s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
Fix $x\in G^{K}$. Since we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$, then the index $n$ with $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite. This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$. Then, by $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon).$ (3.1)
For any $\tau>0$, there exists $\delta_{0}>0$ such that
$\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
For any $k\geq 1$, by definitions, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
$h_{top}^{UC}(f,\mathcal{R}_{K}(\delta_{0},k),\epsilon)\leq\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
Combining (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
@@
Take a strictly decreasing sequence $\{\delta_{k}\}$ of positive real numbers with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
$s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$. We set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex. In the weak-topology, choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed. By Lemma 3.1, we have
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
This implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
Thus, by the Lemma 3.2 the upper bound of the packing metric mean dimension of $G_{K}$ is given by
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.2)
@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
For any small enough $\epsilon^{\ast}>0$, we prove that
@@
Then the above inequality yields the desired lower bound.
Without loss of generality, one may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$ and then take $\mu^{\ast}\in K$ such that
@@
Since $K$ is a connected and compact set, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
@@
Now endow the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with the lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively.
Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is chosen as in the definition of specification property.
@@
For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$, which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2. By the specification property, there exists $a_{k}\in X$ such that
@@
Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$ and $N_{(k,0)}=1$.
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation of $\mu^{*}$ is a slight modification of [LL24, Lemma 2.1], so we omit the proof.
@@
For each $k\in\mathbb{N}$, there exists $\nu_{k}\in M_{f}(X)$ satisfying
@@
$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta;$ (3.5)
$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$, where $D$ is the compatible metric on $M(X)$ given in (2.1).
Since $\nu_{i}^{k}$ is ergodic, by Birkhoff ergodic theorem there exists $N\in\mathbb{N}$ such that the set
@@
has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3, one has
@@
$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer for all $1\leq i\leq j_{k}$, and
@@
$(2)$ there exists a subset $E_{k,i}$ that is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
where the second inequality is derived from Lemma 2.3.
@@
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniformly continuous.
For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$, then
@@
Then $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set and
#$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
@@
${\geq}\exp\left[n_{k}(1-\zeta)\left(H-3\zeta\right)\right],$ (3.7)
where the last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and (3.4), (3.6).
For $n\in\mathbb{N}$ and $\varphi\in C(X)$, we write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$ as the $n$-th Birkhoff sum of $\varphi$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X)$ taking values in $[0,1]$,
@@
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Since
$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k},$
one has that for any $y\in\mathcal{C}_{(k,Q_{k})}$,
@@
Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the cases for $1\leq j<Q_{k}$.
Since the system has specification property, the generic point $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$ [PS07]. Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
@@
With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready for the construction:
Step $1$. Construct the intermediate sets $\mathcal{S}_{(k,j)}$.
@@
For $(k,j)\in\mathcal{A}$ with $j=0$, we set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.